OpenStudy (anonymous):

let u= (3.,4) find two unit vectors that are perpendicular to u

OpenStudy (amistre64):

if u is a vector defined by <3,4>, then we can find 2 vector that are perpendicular to it by the dot product.

OpenStudy (amistre64):

<3,4> . <x,y> = 0 3x +4y = 0

OpenStudy (amistre64):

this is the equation of a line that should be perpendicular to "u". the slope of this line is -3/4 which is good :)

OpenStudy (amistre64):

magnitude doesnt matter so we can pick any vectors pointing in opposite direction on this line

OpenStudy (amistre64):

<-4,3> and <4,-3> should work

OpenStudy (amistre64):

lets test it by the dot ;) <3,4>.<-4,3> = -12 + 12 = 0, thats good <3,4>.<4,-3> = 12 + -12 = 0...good again

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)